OpenStudy (anonymous):

solve the equation and check your solution 10g^2+10=29g

OpenStudy (anonymous):

do you know how to factor quadratic equations?

OpenStudy (anonymous):

no im as far as 10g^2-29g+10=0

OpenStudy (anonymous):

okay. so you do not know how to factor quadratic equations at all. Am I right?

OpenStudy (anonymous):

you can take the coefficients (number multipliers) of the g-squared term and the regular term (which are both ten) and multiply them together to get 100. Then list all the ways you can multiply two numbers together to get 100.

OpenStudy (anonymous):

...EX: 1 * 100 2 * 50 4 * 25 5 * 20 10 * 10

OpenStudy (anonymous):

http://www.youtube.com/watch?v=1Pva-Iv43Nc go through that. and come back here if you still don't get the answer

OpenStudy (anonymous):

then look for one of the pairs that add or subtract to get the middle coefficient: -29 (100 can also come from multiplying two negatives together)

OpenStudy (anonymous):

-4 and -25

OpenStudy (anonymous):

...So once you see that the two factors of 100 that add together to get -29 are -4 and -25, then you re-write the quadratic as: 10g^2 - 25g - 4g + 10 = 0 And work to factor the left side: 5g(2g - 5) - 2(2g - 5) = 0 ...

OpenStudy (anonymous):

...then you continue to factor: (2g - 5)(5g - 2) = 0 and the solution to what makes either parenthetical part equal zero is a solution. The two solutions are solutions to: 2g - 5 = 0 and 5g - 2 = 0

OpenStudy (anonymous):

so g= 5/2 and 2/5 ??

OpenStudy (anonymous):

Yes, personally, I would go into the quadratic formula right away, but that is because I have memorized it, and I am fast at it, while not everybody has or is. Typically it (the QF) is annoying. http://en.wikipedia.org/wiki/Quadratic_formula#Quadratic_formula

OpenStudy (anonymous):

not allowed to use quadratic formula until im a junior in high school and im only a freshmen

OpenStudy (anonymous):

Yes, quadratic formula is sometimes less efficient and less elegant. but for problems with no obvious root, it is the only way

OpenStudy (toxicsugar22):

tonks can u help me

OpenStudy (toxicsugar22):

on my math problem